### v1

@@ 4 The proposed MPGDA algorithm
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\bar{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x})),y-\bar{y}_{k}(\bar{x})\rangle$
$+\langle\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\bar{y}_{k}(\bar{x}),y-\bar{y}_{k}(\bar{x})\rangle.$
Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$,by Cauchy-Schwarz inequality, we arrive at
@@
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
@@
Next, we prove that the descent condition (18) must be satisfied after a unified finite number of iterations for all $k\in N$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
@@
Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in N$ and $0<\alpha\leq\min\{1,~{}\bar{\alpha}\}$, it holds that
@@
By (chenProximalGradientMethod2020, , Lemma 5.2), for any $0<\alpha\leq\min\{1,~{}\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$,
$f(x_{\alpha},\bar{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\bar{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2},$
where $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$ and $G=\sup_{x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\|\nabla_{x}F(x,y)\|$.
@@
$f(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-g(\bar{y}_{k}(x_{\alpha}))$
$\leq f(x_{\alpha},\bar{y}_{k}(x_{k,i}))-g(\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}+\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
By summing these inequalities with the inequality
@@
in Proposition 1, we derive the following inequality:
$F(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$\leq F(x_{k,i},\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, we have
$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$
Note that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$, then for any $0<\alpha\leq\min\{1,\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$, we have
$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
for all $k\geq 0$. ∎
We are now ready to establish the iteration complexity for MPGDA algorithm. Let $\varepsilon>0$ be a given target accuracy, we define
$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k,0}}(x_{k,0},\bar{y}_{k}(x_{k,0}))<\varepsilon\}$
where $\mathcal{G}^{\beta}(\cdot,\cdot)$ is defined in (12). In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for MPGDA algorithm to obtain a $\varepsilon$-game-stationary point. We further define two constants $\overline{Q}$ and $\underline{Q}$ as
$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\},$
$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S\}.$
@@
Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be generated by Algorithm 1. Then, there exists a positive $\bar{\alpha}>0$ such that the following inequality holds:
$T(\varepsilon)\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{\varepsilon^{3}}$
@@
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}}.$
Here, all the parameters $\rho_{0},~{}\gamma_{0},~{}\sigma_{y},~{}\overline{Q},~{}\underline{Q},~{}\overline{T},~{}S_{\rho},~{}l_{\max},~{}\eta$ are defined as in the relevant context of the paper.
@@
Let $\beta_{k}:=\beta_{k,0},~{}v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$ and $\rho_{k+1}\leq\rho_{k}$, by the definition of $Q_{k}$, we have
$Q_{k+1}(x_{k+1})$ $=h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
$\leq h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+4\rho_{k}\sigma_{y}^{2}$
$=Q_{k}(x_{k,T_{k}})+4\rho_{k}\sigma_{y}^{2}.$
Furthermore, according to Proposition 5, there exists $\bar{\alpha}>0$ such that
$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+2)\rho_{k}\sigma_{y}^{2}-\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}.$
@@
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K})+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}$ (20)
with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\geq(\frac{\bar{\alpha}\eta\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}k^{1/3})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (21)

### v2

@@ 4 The proposed MPGDA-PA algorithm
$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\overline{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\overline{y}_{k}(\bar{x})),y-\overline{y}_{k}(\bar{x})\rangle$
$+\langle\rho_{k}(\overline{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\overline{y}_{k}(\bar{x}),y-\overline{y}_{k}(\bar{x})\rangle.$
Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$ and $\nabla_{y}f(x,y)$ is $L_{y}$-Lipschitz continuous, it follows that,
@@
$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
@@
Next, we prove that the descent condition (24) must be satisfied after a unified finite number of iterations for all $k\in\mathbb{N}$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
@@
Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in\mathbb{N}$ and $0<\alpha\leq\min\{1,~\bar{\alpha}\}$, it holds that
@@
Let $G:=\sup\{\|\nabla_{x}f(x,y)\|\mid x\in\mathcal{M},y\in S\}<\infty.$ Assume that $0<\alpha\leq\min\{1,~\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$ with $M_{1}>0,~M_{2}>0$ being defined in Proposition 1 and $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$. Then, following an argument similar to the proof of (chenProximalGradientMethod2020, Lemma 5.2), we deduce that
$f(x_{\alpha},\overline{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\overline{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2}.$
@@
$f(x_{\alpha},\overline{y}_{k}(x_{\alpha}))-g(\overline{y}_{k}(x_{\alpha}))$
$\leq f(x_{\alpha},\overline{y}_{k}(x_{k,i}))-g(\overline{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{k,i})\|^{2}$
$+\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{\alpha})\|^{2}+\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
By summing the above two inequalities and the inequality
@@
from Proposition 1, we derive the following inequality:
$F(x_{\alpha},\overline{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{\alpha})\|^{2}$ (26)
$\leq F(x_{k,i},\overline{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{k,i})\|^{2}$
$+\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, it holds that
$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$ (27)
Combining (LABEL:ineq-F-gammak), (27), and the definition of $Q_{k}(x)$, we obtain that for all $k\geq 0$ and any $0<\alpha\leq\min\{1,\bar{\alpha}\}$,
$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2},$
where $\bar{\alpha}=\min\{\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$. ∎
By Proposition 5, the backtracking line search procedure in Algorithm 1 will terminate after a finite number of steps. We present this result in the following proposition.
Proposition 6.
For all $k\geq 0$, the condition (24) in Algorithm 1 is satisfied within at most $J_{1}$ backtracking steps, where
[…]
We are now ready to establish the iteration complexity for Algorithm 1. Let $\varepsilon>0$ be a given target accuracy, we define
$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k}}(x_{k},y_{k})<\varepsilon\}$ (28)
where $\mathcal{G}^{\beta}(\cdot,\cdot)$ is defined in (13). In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for Algorithm 1 to obtain a $\varepsilon$-game-stationary point. We further introduce two constants $\overline{Q}$ and $\underline{Q}$ as follows
$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in S}\},$
$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in S\}.$
@@
Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be the sequence generated by Algorithm 1. Then, the following inequality holds:
$T(\varepsilon)\leq\frac{(\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\})^{3}}{\varepsilon^{3}}+1$
@@
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+\frac{\gamma_{0}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3c_{1}\eta^{J_{1}}\gamma_{0}}}.$
Here, the constants $c_{1},~\eta,~\rho_{0},~\gamma_{0},~\overline{T}$, and $l_{\max}$ are parameters of Algorithm 1, $J_{1}$ is defined in Proposition 6, and $S_{\rho}:=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
@@
Let $\beta_{k}:=\beta_{k,0},~v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$, $\rho_{k+1}\leq\rho_{k}$, and $x_{k+1}=x_{k,T_{k}}$, by the definition of $Q_{k}$ and $\sigma_{y}$, we have
$Q_{k+1}(x_{k+1})$
$=$ $h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
$\leq$ $h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}$
$+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}+2\rho_{k}\sigma_{y}^{2}$
$=$ $Q_{k}(x_{k,T_{k}})+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}+2\rho_{k}\sigma_{y}^{2}.$
Combining this with (24) and Proposition 6, we obtain that
$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+1)\rho_{k}\sigma_{y}^{2}+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}-c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}.$
@@
$\sum_{k=K_{1}}^{K}c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K+1})+\frac{\gamma_{K_{1}}-\gamma_{K}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2}$ (29)
with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$. Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
$\sum_{k=K_{1}}^{K}c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}\geq(\frac{c_{1}\eta^{J_{1}}\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}\frac{1}{k^{1/3}})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (30)